A fruit processor manufactures and sales a single product.the company's total sales, variable cost and fixed cost are $ 60000, $420000 and $150000 respectively. if the unit sellin\sqrt(g) price is $40. calculate the breakeven point the unit sold and total revenu.


1
Expert's answer
2021-11-09T01:02:45-0500

a.

Break-Even point (units) = Fixed Costs "\div" (Sales price per unit - Variable costs per unit)


variable cost per unit is = "\frac{\$420000}{no. of units}"


Units= "\frac{total sales}{selling price}"


="\frac{\$600000}{40}"


=15000 units sold


so now the variable cost per unit is= "\frac{\$420000}{15000}=\$28"


breakeven in units


= "\frac{\$150000}{\$(40-28)}=12,500" units


Total revenue:

=selling price per unit "\times" total units sold


"=\$40 \times15000=\$600,000"